# A Survey of Generative AI for Intelligent Transportation Systems

### v1

@@ 4.1. Human-related Traffic Prediction
Individual behavior. We discuss individual behavior prediction based on their roles within transportation systems. For drivers, modeling and predicting driver behavior plays a crucial role in enhancing safety and preventing risky actions (Ivanovic et al., 2020; Bao et al., 2021). For example, Bao et al. (Bao et al., 2021) introduced a probabilistic sequence-to-sequence approach utilizing a conditional VAE to predict various driving behaviors. For pedestrians, accurately forecasting human motion behavior is vital for autonomous vehicles to plan preemptive actions and avoid collisions, ensuring pedestrian safety (Sun et al., 2019; Cui et al., 2021). For instance, in (Cui et al., 2021), the authors proposed a temporal convolutional GAN for high-fidelity future pose prediction, incorporating two discriminators: one for fidelity assessment and another for consistency evaluation in long sequences.
Human mobility. Human mobility refers to an individual’s movement from one location to another. Feng et al. (Feng et al., 2023b) employed GAIL to model courier decision-making in route choice, enabling the prediction of their future routes. The COVID-19 pandemic presented a great challenge in predicting human mobility. Bao et al. (Bao et al., 2020) introduced a spatio-temporal conditional GAN Covid-GAN to estimate mobility during the pandemic. Further, Covid-GAN+ (Bao et al., 2022) addressed spatial heterogeneity by using the local Moran statistic and outliers by redesigning the training objective to learn the estimated mobility changes.
@@ 4.2. Vehicle-related Traffic Prediction
Vehicle trajectory. A vehicle trajectory refers to the path a vehicle takes from one location to another over a specific time period. Precise vehicle trajectory prediction is vital for a variety of traffic-related applications, including traffic management, navigation, and route planning. However, achieving precise trajectory predictions is a challenging task for several reasons. First, vehicle movements are influenced by various factors, including driver behavior, traffic conditions, road infrastructure, and unexpected events. Second, traffic conditions can change rapidly due to incidents, congestion, or accidents. Third, the accuracy of trajectory prediction models heavily relies on the quality and quantity of input data, such as GPS data. Noisy or incomplete data can lead to inaccurate predictions. To address these challenges, more researchers proposed different methods based on generative AI technology to predict vehicle trajectory.
@@
Most problems of vehicle trajectory prediction are defined as using historical trajectories to infer future trajectories. Formally, given a vehicle $i$ with a historical trajectory $\mathbf{x}_{i}=\{r_{i}^{t}\},t=1,2,...,T$, where $r_{i}^{t}$ is location information (e.g., longitude and latitude) at time $t$. The task is to learn a mapping function $g(\cdot)$ to predict the future trajectory $\mathbf{y}_{i}=\{r_{i}^{t}\},t=T+1,...,T+n$. Many scholars adopt different generative AI techniques to address above challenges.
@@
VAE. Several researchers have designed prediction models based on VAE (Cheng et al., 2020; Kim et al., 2021; Xing et al., 2022; Oh and Peng, 2022). DVAE (Neumeier et al., 2021) incorporated expert knowledge into the decoder of a VAE, which can provide the similar prediction accuracy and enhance interpretability. The authors in (Jagadish et al., 2021, 2022) utilized conditional VAEs to encode past trajectories of vehicles and traffic scenes for vehicle path prediction. Gui et al. (Gui et al., 2022) incorporated more information like yaw angle and velocity in their proposed conditional VAE-based trajectory prediction model.
GAN. GAN is also an alternative technique for trajectory prediction (Roy et al., 2019; Li et al., 2021a; Zhou et al., 2021b). For example, Kang et al. (Kang et al., 2020a) applied social GAN, originally designed for pedestrian trajectory prediction, for the prediction of vehicle trajectories. Differently, Zhou et al. (Zhou et al., 2021b) focused on predicting trajectories of surrounding vehicles. They introduced a self-attention social GAN, which not only addressed the challenge of missing information from long trajectory sequences but also better modeled the interactions between vehicles.
@@
GPT. Recently, GPT models has great potential in vehicle trajectory prediction (Su et al., 2022; Feng et al., 2023a; Vishnu et al., 2023). Su et al. (Su et al., 2022) proposed a crossmodal transformer based generative framework, which leveraged sequences of cues from various modalities and pedestrian attributes for more accurate predictions. In this model, transformer is adopted to capture the cross-relation patterns of different modality pairs. Feng et al. (Feng et al., 2023a) developed a pre-trained large traffic model using the Transformer architecture. This model has ability to capture the diverse trajectories within a population of vehicles, leading to higher accuracy in vehicle trajectory prediction.
@@
Crash risk. Crash risk prediction is important in traffic safety management and precaution. Due to imbalanced data between crash and non-crash cases, accurately predicting crash risk is challenging. To deal with this, Cai et al. (Cai et al., 2020) proposed a deep convolutional GAN model to generate more synthetic data about crashes. Similarly, in (Man et al., 2022), Wasserstein GAN was developed to address the imbalanced data. Different from approaches using GAN, Zhang et al. (Zhang et al., 2022d) proposed an enhanced accelerated testing method based on Normalizing Flows. This method leverages Normalizing Flows to learn the distribution of rare events and preserve the correlated relationships between scenario variables.
@@ 4.3. Road Segment-related Traffic Prediction
VAE. Some researches propose VAE-based models. The authors in (Boquet et al., 2020) designed a VAE-based prediction model with the ability to learn the data distribution. PFVAE (Jin et al., 2022a) employed LSTM as the autoencoder and designed the VAE as a predictor for time-series data to deal with the noise issues.
@@
GAN. GAN-based models are commonly adopted in this task (Lin et al., 2018; Zhang et al., 2019b; Saxena and Cao, 2019; Zhang et al., 2021e; Devadhas Sujakumari and Dassan, 2023). For example, ForGAN (Koochali et al., 2019) employed GAN to learn the data generating distribution and generate probabilistic traffic forecasts. Furthermore, many researchers designed several methods to model spatial and temporal correlations in their proposed GAN-based prediction model (Zang et al., 2019; Xu et al., 2020a; Yu et al., 2020; Zheng et al., 2022; Jin et al., 2022b; Khaled et al., 2022; Li et al., 2022b; Zhao et al., 2022b). For example, PL-WGAN (Jin et al., 2022b) utilized GCN, RNN and attention mechanism to model spatiotemporal correlations in a Wasserstein GAN-based model. ASTGAN (Liu and Zhang, 2022) used the attention mechanism and an mask graph convolutional recurrent network to characterize the temporal and spatial dependencies.
@@
Normalizing flows. Consider the capability of conditional normalizing flows in representing complex distributions characterized by high dimensionality and strong interdimensional relationships. MotionFlow (Zand et al., 2023) utilized conditional normalizing flows for traffic prediction, which conditionally models the output distributions on spatio-temporal input features in an autoregressive manner.
@@
DPM. Because DPM can effectively handle uncertainties within spatio-temporal graph neural network (STGNN), DiffSTG (Wen et al., 2023) combined the spatiotemporal learning capability of STGNN with the uncertainty measurement of DPM, resulting in higher prediction accuracy. The authors in (Rasul et al., 2021) developed an autoregressive denoising diffusion model for multivariate probabilistic time series forecasting, utilizing gradient estimation to draw samples from the data distribution at every time step.
@@
Travel time estimation. Travel time estimation task aims to estimate the duration it takes to travel between two locations or along a specific route given the departure time. Several scholars employ generative AI techniques to address this task (Song et al., 2021; Lin et al., 2023b). To be specific, GCGTTE (Song et al., 2021) integrated GCN and GAN to estimate the travel time in a probabilistic distribution form. Lin et al. (Lin et al., 2023b) proposed a two-stage diffusion-based origin-destination (OD) travel time estimation model. They incorporated a conditioned pixelated trajectory denoiser into the model to capture correlations between OD pairs and historical trajectories.
@@ 4.4. Region-specific Traffic Prediction
Region-specific traffic prediction focuses on forecasting traffic conditions such as traffic flow within specific geographic regions in a transportation network. The main challenges in achieving accurate predictions include complex spatio-temporal dependencies, limited data, and noise. Recent works (Li et al., 2020a; Zhang et al., 2020c, 2019a; Naji et al., 2021; Zhang et al., 2023a; Feng et al., 2021) have aimed to employ generative AI techniques to tackle these challenges. To be specific, Zhang et al. (Zhang et al., 2022b, 2023a) proposed a spatiotemporal GAN model for predicting traffic flow in subway stations, which used gated TCN and weight sharing GCN to capture spatiotemporal correlations. Kang et al. (Kang et al., 2020b) incorporated gated convolution and dilated convolution to capture the traffic patterns in urban regions for region-level traffic prediction. Li et al. (Li et al., 2022a) proposed a seq2seq spatial-temporal semantic GAN model, which takes into account both semantic factors and spatio-temporal features. The authors in (Naji et al., 2021) incorporated recurrent network mode and conventional network model into GAN to predict the taxi demand in specific areas.
@@ 5.1. Driver Behavior Simulation
In autonomous driving scenarios, the focus is on how drivers make decisions in response to traffic conditions (Ghosh et al., 2016; Yun et al., 2019; Chen et al., 2022b; Ye et al., 2022; Singh et al., 2023). Yun et al. (Yun et al., 2019) utilized GAN to learn the human-driving abilities in complex traffic scenarios. Chen et al. (Chen et al., 2022b) proposed a novel two-stage driver model that integrates model-based and data-driven techniques, leveraging reward-augmented GAIL to produce reliable human driving behaviors. Jin et al. (Jin et al., 2023) proposed a generative driver agent simulation framework built upon LLM to simulate human driving behaviors based on human driver experience. To alleviate the computational load of large-scale simulations, Ye et al. (Ye et al., 2022) proposed an efficient method for dynamically calculating agent decision parameters in ABM calibration by using VAEs to accelerate state transfer probability computation.
Lane changing and car following are two common driving behaviors, and numerous studies have explored methods for simulating these behaviors (Lin et al., 2021; Ma et al., 2023; Li et al., 2023b; Dong et al., 2023). Ma et al. (Ma et al., 2023) introduced a physics-informed conditional GAN that integrates the advantages of both physics-based and deep learning models to to improve the modeling of car-following behavior over multiple time steps in complex traffic flow environments. Lin et al  (Lin et al., 2021) introduced a constrained-GAIL framework that employs reward augmentation to guide driving agents in avoiding undesirable states by imposing additional constraints through a manually designed reward function. In (Dong et al., 2023), the authors leveraged the transformer and diffusion models to generate lane-changing trajectories that closely mimic human behavior.
Unlike autonomous driving scenarios, in transportation services, a variety of factors such as estimated travel demand tend to dominate drivers’ behaviors. For example, Zhang et al. (Zhang et al., 2020a) developed conditional GAIL model capable of learning the driver’s decision-making tendencies and strategies through knowledge transfer among taxi driver agents and across various locations.

### v2

@@ 4.1. Human-related Traffic Prediction
Individual behavior. We discuss individual behavior prediction based on their roles within transportation systems. For drivers, modeling and predicting driver behavior plays a crucial role in enhancing safety and preventing risky actions (Ivanovic et al., 2020; Bao et al., 2021). For example, Bao et al. (Bao et al., 2021) introduced a probabilistic sequence-to-sequence approach utilizing a conditional VAE to predict various driving behaviors. For pedestrians, accurately forecasting human motion behavior is vital for autonomous vehicles to plan preemptive actions and avoid collisions, ensuring pedestrian safety (Sun et al., 2019; Cui et al., 2021). For instance, in (Cui et al., 2021), the authors proposed a temporal convolutional GAN for high-fidelity future pose prediction, incorporating two discriminators: one for fidelity assessment and another for consistency evaluation in long sequences. UTD-PTP (Tang et al., 2024) introduces a universal transformer diffusion modeling framework that combines transformer architectures with diffusion-based uncertainty modeling to accurately predict pedestrian trajectories in complex autonomous driving environments.
Human mobility. Human mobility refers to an individual’s movement from one location to another. Feng et al. (Feng et al., 2023b) presented a graph-based imitation learning framework that adeptly models courier decision-making in route selection. By utilizing a multi-graph neural network, ILRoute effectively captured the complex, multi-source features influencing route choices, while incorporating mobility regularity and personalized constraints to navigate the expansive decision-making space. This approach improved prediction accuracy across multiple real-world datasets and online evaluations. Building on the necessity to understand and predict human mobility patterns during unprecedented events like the pandemic, Bao et al. (Bao et al., 2020) introduced a spatio-temporal conditional GAN Covid-GAN to estimate mobility changes under varying public health conditions and policy interventions. Further, the authors extended their initial model to Covid-GAN+ (Bao et al., 2022) addressed spatial heterogeneity by using the local Moran statistic and outliers by redesigning the training objective to learn the estimated mobility changes. This method achieved a more accurate and robust estimation of human mobility.
@@ 4.2. Vehicle-related Traffic Prediction
Vehicle trajectory. A vehicle trajectory refers to the path a vehicle takes from one location to another over a specific time period. Precise vehicle trajectory prediction is vital for a variety of traffic-related applications, including traffic management, navigation, and route planning.
@@
Most problems of vehicle trajectory prediction are defined as using historical trajectories to infer future trajectories. Formally, given a vehicle $i$ with a historical trajectory $\mathbf{x}_{i}=\{r_{i}^{t}\}$ for $t=1,2,\dots,T$, where $r_{i}^{t}$ represents the location at time $t$, the goal is to learn a mapping function $g(\cdot)$ to predict the future trajectory $\mathbf{y}_{i}=\{r_{i}^{t}\}$ for $t=T+1,\dots,T+n$.
However, achieving precise trajectory predictions is a challenging task for several reasons. First, vehicle movements are influenced by various factors, including driver behavior, traffic conditions, road infrastructure, and unexpected events. Second, traffic conditions can change rapidly due to incidents, congestion, or accidents. Third, the accuracy of trajectory prediction models heavily relies on the quality and quantity of input data, such as GPS data. Noisy or incomplete data can lead to inaccurate predictions. To address these challenges, more researchers proposed different methods based on generative AI technology to predict vehicle trajectory.
@@
VAE. Several researchers have designed prediction models based on VAE (Cheng et al., 2020; Kim et al., 2021; Xing et al., 2022; Oh and Peng, 2022). DVAE (Neumeier et al., 2021) integrated expert knowledge into the decoder component of the VAE, which can provide the similar prediction accuracy and enhance interpretability. The authors in (Jagadish et al., 2021, 2022) utilized conditional VAEs to encode past trajectories of vehicles and traffic scenes for vehicle path prediction. By conditioning the network on these inputs, their approach effectively captured the multimodal nature of vehicle movements in various traffic scenarios. Gui et al. (Gui et al., 2022) incorporated additional dynamic information like yaw angle and velocity in their proposed conditional VAE framework. This approach allows for a more comprehensive understanding of vehicle interactions and movements, resulting in superior performance on both established and newly introduced simulated datasets.
GAN. GAN is also an alternative technique for trajectory prediction (Roy et al., 2019; Li et al., 2021a; Zhou et al., 2021b). For example, Kang et al. (Kang et al., 2020a) applied social GAN, originally designed for pedestrian trajectory prediction, for the prediction of vehicle trajectories. While this approach demonstrated better prediction accuracy, its reliance on models initially tailored for pedestrian dynamics may limit its adaptability to the distinct behaviors exhibited by vehicles. Differently, Zhou et al. (Zhou et al., 2021b) introduced the self-attention social GAN to handle long trajectory sequences and complex vehicle interactions. The self-attention mechanism helped capture correlations in trajectory data, reducing the loss of important information over time. However, the model’s complexity may limit its suitability for real-time applications requiring high computational efficiency. KI-GAN (Wei et al., 2024) was a knowledge-informed GAN that integrated traffic signal data and multi-vehicle interactions through a specialized attention pooling mechanism. This model effectively incorporated orientation and proximity information for accurate predictions but relied heavily on the quality of input data, which can vary across urban environments.
@@
GPT. Recently, GPT models has great potential in vehicle trajectory prediction (Su et al., 2022; Feng et al., 2023a; Vishnu et al., 2023). Su et al. (Su et al., 2022) proposed a crossmodal transformer based generative framework that integrates multiple modalities and pedestrian attributes to predict pedestrian trajectories more accurately. Similarly, Feng et al. (Feng et al., 2023a) developed a pre-trained large traffic model using the Transformer architecture. This model has ability to capture the diverse trajectories within a population of vehicles, leading to higher accuracy in vehicle trajectory prediction. However, the high computational demands of such large model may pose challenges for real-time deployment in autonomous systems.
@@
Crash risk. Crash risk prediction is important in traffic safety management and precaution. Due to imbalanced data between crash and non-crash cases, accurately predicting crash risk is challenging. To deal with this, Cai et al. (Cai et al., 2020) proposed a deep convolutional GAN model to generate synthetic crash data. This approach effectively preserves the heterogeneity of non-crash scenarios, enabling the use of the entire non-crash dataset for training predictive models. Similarly, in (Man et al., 2022), Wasserstein GAN (WGAN) was developed to address the extreme class imbalance. By generating realistic synthetic crash instances, WGAN achieved higher sensitivity in crash detection while maintaining a low false alarm rate. Different from approaches using GAN, Zhang et al. (Zhang et al., 2022d) proposed an enhanced accelerated testing method based on Normalizing Flows. This method efficiently learned the distribution of rare crash events and preserves the correlations between scenario variables, significantly reducing the number of required tests by over two hundred times compared to traditional Monte Carlo simulations. However, the method’s effectiveness is highly dependent on the quality of the learned distributions and may require extensive computational resources for training the Normalizing Flows models.
@@ 4.3. Road Segment-related Traffic Prediction
VAE. Some researches have proposed VAE-based models. The authors in (Boquet et al., 2020) designed a VAE-based prediction model with the ability to learn the data distribution, resulting in improved forecasting performance and versatility across multiple real-world datasets. PFVAE (Jin et al., 2022a) presented a novel planar flow-based VAE for traffic flow prediction. By integrating LSTM as the autoencoder and employing planar flows within the VAE structure, the model effectively captured the intricate temporal dependencies and nonlinear patterns of time-series data. Empirical results demonstrated its robustness and accuracy in handling noisy time-series data.
@@
GAN. GAN-based models are commonly adopted in this task (Lin et al., 2018; Zhang et al., 2019b; Saxena and Cao, 2019; Zhang et al., 2021d; Devadhas Sujakumari and Dassan, 2023). For example, ForGAN (Koochali et al., 2019) employed GAN to learn the underlying data distribution for accurate probabilistic forecasts. This approach not only enhanced prediction accuracy but also provided a robust framework for handling real-world fluctuations in time-series data. In particular, many researchers designed several methods to model spatiotemporal correlations in GAN-based prediction model (Zang et al., 2019; Xu et al., 2020a; Yu et al., 2020; Zheng et al., 2022; Jin et al., 2022b; Khaled et al., 2022; Li et al., 2022b; Zhao et al., 2022b). For example, PL-WGAN (Jin et al., 2022b) utilized GCN, RNN and attention mechanisms within a Wasserstein GAN framework to model spatiotemporal correlations in urban traffic networks, leading to highly accurate short-term traffic speed predictions. However, the combination of multiple neural network components increased computational demands. Similarly, ASTGAN (Liu and Zhang, 2022) used the attention mechanism and an mask graph convolutional recurrent network to characterize temporal and spatial dependencies, improving data completeness and prediction accuracy. The GAN component ensured the reliability of forecasts, showing significant improvements in real-world traffic prediction tasks.
@@
Normalizing flows. Consider the capability of conditional normalizing flows in representing complex distributions characterized by high dimensionality and strong interdimensional relationships. MotionFlow (Zand et al., 2023) utilized conditional normalizing flows to model complex spatio-temporal data distributions. This model effectively captured the intricate variability inherent in traffic patterns through its autoregressive conditioning on spatio-temporal inputs.
@@
DPM. Since DPM can effectively handle uncertainties within spatio-temporal graph neural network (STGNN), DiffSTG (Wen et al., 2023) combined the spatiotemporal learning capability of STGNN with the uncertainty measurement of DPM, resulting in higher prediction accuracy. The authors in (Rasul et al., 2021) developed an autoregressive denoising diffusion model for multivariate probabilistic time series forecasting, utilizing gradient estimation and Langevin sampling to accurately sample data distribution. Experimental results showed that it outperformed traditional and contemporary forecasting models in both accuracy and robustness.
@@
Travel time estimation. Travel time estimation task aims to estimate the duration it takes to travel between two locations or along a specific route given the departure time. Several scholars employ generative AI techniques to address this task (Song et al., 2021; Lin et al., 2023b). To be specific, GCGTTE (Song et al., 2021) integrated GCN and GAN to estimate travel time in a probabilistic form. This approach not only enhances prediction accuracy but also provides valuable insights into traffic variability, which is essential for proactive traffic management. Lin et al. (Lin et al., 2023b) proposed a two-stage diffusion-based origin-destination (OD) travel time estimation model. The model incorporated a conditioned pixelated trajectory denoiser to capture correlations between OD pairs and their historical trajectories, effectively managing variability and noise in trajectory data. It outperformed baseline methods in terms of accuracy, scalability, and explainability.
@@ 4.4. Region-specific Traffic Prediction
Region-specific traffic prediction focuses on forecasting traffic conditions such as traffic flow within specific geographic regions in a transportation network. The main challenges in achieving accurate predictions include complex spatio-temporal dependencies, limited data, and noise. Recent works (Li et al., 2020a; Zhang et al., 2020c, 2019a; Naji et al., 2021; Zhang et al., 2023a; Feng et al., 2021) have aimed to employ generative AI techniques to tackle these challenges. To be specific, Kang et al. (Kang et al., 2020b) developed a GAN-based model that integrates gated and dilated convolutions to effectively capture both local and distant traffic patterns within urban regions. This dual convolution approach allows for more precise real-time traffic flow predictions compared to traditional grid-based models. Li et al. (Li et al., 2022a) proposed a seq2seq spatial-temporal semantic GAN model for multi-step urban flow prediction. By treating successive urban flow data as video frames, the model employed a spatial-temporal semantic encoder to simultaneously capture semantic factors and spatial-temporal dependencies. The integration of adversarial loss with prediction error effectively mitigates the issue of blurry predictions, resulting in more accurate and reliable forecasts. The authors in (Naji et al., 2021) incorporated recurrent and conventional network models into the GAN framework to predict the taxi demand in specific areas. The model demonstrated superior performance on real-world data, and provided valuable insights into the dynamics of taxi supply and demand.
4.5. Discussion
Generative AI has made great progress in traffic prediction, addressing challenges in data quality, behavioral modeling, and spatiotemporal dynamics. Despite their high accuracy, these models often lack interpretability, which is crucial for building trust in ITS. Approaches like DVAE (Neumeier et al., 2021), which incorporate expert knowledge, improve interpretability without sacrificing performance. This balance is crucial for gaining trust and facilitating the adoption of generative AI technology in ITS. Real-time applications also present computational challenges. While GPT-based frameworks (Su et al., 2022; Feng et al., 2023a; Vishnu et al., 2023) offer high prediction accuracy, they struggle with latency and resource demands, highlighting the need for optimized models that balance performance and efficiency. In summary, achieving a balance between accuracy, interpretability, and efficiency remains critical in traffic prediction. Future research should focus on developing generative AI models that not only achieve high accuracy but also maintain transparency and efficiency in real-time traffic environments.
@@ 5.1. Driver Behavior Simulation
In autonomous driving scenarios, the focus is on how drivers make decisions in response to traffic conditions (Ghosh et al., 2016; Yun et al., 2019; Chen et al., 2022b; Ye et al., 2022; Singh et al., 2023). Yun et al. (Yun et al., 2019) utilized GAN to develop a data-driven, human-like driver model capable of simulating intricate interactions in traffic scenarios. By leveraging GANs, their model effectively captured the intricate driving skills exhibited by humans in diverse and complex situations, outperforming traditional rule-based models. Chen et al. (Chen et al., 2022b) introduced a two-stage hybrid driver model that integrates data-driven neural networks with model-based controllers. By employing reward-augmented GAIL, their model effectively simulated human-like driving behaviors while ensuring safety and expressiveness in multi-agent highway scenarios. In a novel application of LLM, Jin et al. (Jin et al., 2023) developed a generative driver agent simulation framework that leverages human driver experience through self-reported driving thoughts. By integrating LLMs with human-derived chain-of-thought data, their framework significantly reduces collision rates and enhances the human-like qualities of simulated drivers. Addressing the computational challenges associated with large-scale simulations, Ye et al. (Ye et al., 2022) proposed an efficient calibration method for agent-based traffic simulations using VAEs. By compressing agent state vectors into lower-dimensional latent spaces, their approach dramatically reduced the computational burden of calculating state transfer probabilities.
Lane changing and car following are two common driving behaviors, and numerous studies have explored methods for simulating these behaviors (Lin et al., 2021; Ma et al., 2023; Li et al., 2023b; Dong et al., 2023). Ma et al. (Ma et al., 2023) introduced a physics-informed conditional GAN that integrates the advantages of both physics-based models and deep learning techniques. This hybrid approach eliminated the need for explicit weighting parameters, simplifying the model while effectively capturing multi-step car-following behaviors in complex traffic environments. Lin et al  (Lin et al., 2021) introduced a constrained-GAIL framework that employs reward augmentation to guide driving agents away from undesirable behaviors. By incorporating manually designed reward functions alongside those learned from data, their model achieved higher prediction accuracy in vehicle speed and location while significantly reducing collisions and erratic driving behaviors. Addressing the need for highly realistic simulation environments, the authors in (Dong et al., 2023) leveraged the transformer and diffusion models to generate human-like lane-changing trajectories. Extensive experiments validated its ability to produce realistic and controllable trajectories.
Unlike autonomous driving scenarios, in transportation services, a variety of factors such as estimated travel demand tend to dominate drivers’ behaviors. For example, Zhang et al. (Zhang et al., 2020a) developed conditional GAIL model capable of learning the driver’s decision-making tendencies and strategies through knowledge transfer across different drivers and locations. By integrating collective inverse RL, cGAIL enhances the generalization of driver behaviors, thereby improving traffic management and service quality.